OpenStudy (anonymous):

write an equation that describes a line that passes through (7,1) and is perpendicular to the line described by y=-1/2x+3

OpenStudy (jhannybean):

So you need to use the point slope form to solve this equation: y-y1=m(x-x1) where the slope is m=-(-2), the negative reciprocal of the original because the equation you're trying to find is perpendicular to the line given

OpenStudy (jhannybean):

Using the point slope form with m=2, and point (7,1) plug in your values. Y-y1=m(x-x1) Y-1=2(x-7) Y=2x-13